OpenStudy (anonymous):

Simplify 625 to the negative 1 fourth power

OpenStudy (unklerhaukus):

\[625^{-1/4}=\frac1{625^{1/4}}=\frac1{\sqrt{625}^{1/2}}=\frac1{\sqrt{25}}=\frac{1}{5}\]